# On lower bounds for the distances between APN functions

### v1

@@ 4.2 Properties of the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is plateaued APN
Our lower bound in Theorem 4.13 is a dramatic increase over the previously known $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 1$ when $F$ is plateaued APN. It would be interesting to further improve this lower bound or find an example of a family of plateaued APN functions for which it is tight. The lower bound on the distance from plateaued APN functions to other APN functions immediately follows.
@@
When $n$ is odd, any plateaued APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is AB, and so we know by a result of [35] that $\mathcal{G}_{F}$ is a maximal Sidon set with all points outside of the graph having exclude multiplicity $\frac{2^{n}-2}{6}$ which is odd (see Section 4.1). However, when $n$ is even, the exclude multiplicities of $\mathcal{G}_{F}$ are much more unclear. Computational results seem to suggest that if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued APN function, then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and we conjecture in 4.15 that this is true. In the following proposition, we describe the parity of the exclude multiplicities of $\mathcal{G}_{F}$ for $F$ plateaued APN in terms of the Fourier transform of the indicator function $1_{\mathcal{B}({F})}\colon\mathbb{F}_{2}^{n}\to\left\{0,1\right\}$ of $\mathcal{B}({F})$}.
@@
Assume $n$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function and let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(F(a)+b)\equiv 2\mod 4$. Equivalently, $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even.
@@
For all $v\in\mathbb{F}_{2}^{n}$, let $\lambda_{v}$ denote the amplitude of $v\cdot F$, and let $m_{v}$ be the non-negative integer such that $\lambda_{v}^{2}=2^{n+m_{v}}$. Note that $m_{v}$ is even and $0\leq m_{v}\leq n$ for all $v\in\mathbb{F}_{2}^{n}$. Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and let $c=F(a)+b$. By Proposition 4.7, we have
@@
So $6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\equiv\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\mod 12$. Note that $\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 6\mod 12$ if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ and $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 0\mod 3$. However, we already know that $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)=0$ as $c\neq 0$. So $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$, and this proves the first part of the statement.
For $i\in\left\{0,1\right\}$, let $B_{i}=|\left\{v\in{\mathcal{B}({F})}:v\cdot c=i\right\}|$. Then $\widehat{1_{\mathcal{B}({F})}}(c)=|B_{0}|-|B_{1}|=|{\mathcal{B}({F})}|-2|B_{1}|$. It was shown in [30, Proposition 6.5] that the number of bent components of any plateaued function over an even dimension is $2\mod 4$, and so $|{\mathcal{B}({F})}|\equiv 2\mod 4$. Therefore, $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2-2|B_{1}|\mod 4$. So $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ if and only if $|B_{1}|$ is even. Since the parity of $|B_{0}|$ and $|B_{1}|$ is the same, the second part of the result follows. ∎
Conjecture 4.15.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
In Section 4.4, we prove the above conjecture in the affirmative in the case of $F$ being quadratic APN, but we were unable to find a proof for all plateaued APN functions.
We now express the exclude multiplicities of the graph of an APN function in terms of $\widehat{W_{F}^{4}}$.
@@
$=|\left\{(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(y)+F(z)+F(x+y+z)=(0,c)\right\}|$
@@
From the above result, we can also express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of the derivatives of $\gamma_{F}$. In Section 7, we will see why this is relevant and how it is connected to the existence of nontrivial linear structures of $\gamma_{F}$. In [13, Section 6], it was shown that for any APN function $F$, we have $W_{F}^{2}=2^{2n}\delta_{(0,0)}-W_{\gamma_{F}}+2^{n}$, (where $\delta_{(0,0)}$ takes value $1$ at $(0,0)$ and 0 elsewhere). So
@@
Throughout this subsection, we assume that $n$ is even. We say a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is 33-to-11 if every point in the image of $F$ has a preimage of size $3$ except a single point. Note that all 3-to-1 plateaued functions are necessarily APN (see [29, Theorem 12]).
@@
Moreover, [7, Corollary 4] describes the multiplicities of the nonzero elements of the multiset $\left\{F(x)+F(y)+F(x+y):x,y\in\mathbb{F}_{2}^{n}\right\}$ when $F$ is quadratic 33-to-11, and indeed, this can be easily generalized to all plateaued 3-to-1 APN functions.
@@
These two results above (in terms of quadratic 3-to-1 functions) and [5, Proposition 5] (which states that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$ when $F$ is quadratic) were used in [7] to determine the exact exclude multiplicities of $\mathcal{G}_{F}$ and their frequencies in the case of $F$ being a quadratic 33-to-11 function. This can be done because any plateaued function $F$ with all component functions unbalanced satisfies
@@
for all $x,w\in\mathbb{F}_{2}^{n}$ (see [10, Theorem 2]). Therefore, if $F$ is plateaued APN with all unbalanced components and $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, then
@@
Without loss of generality, assume $F(0)=0$. By Theorem 4.18, $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$. The result then immediately follows by applying Proposition 4.4 and Corollary 4.19 (cf. [7]) and using the same proof as [7, Corollary 4]. ∎
@@ 4.4 Quadratic APN functions
$\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F})).$ (9)
@@
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$. For any $b\neq 0$, the weight of the component function $b\cdot\pi_{F}$ of the ortho-derivative of $F$ satisfies $\operatorname{wt}(b\cdot\pi_{F})\leq 2^{n}+2-3\cdot 2^{\frac{n}{2}-1}$.
@@
It was also shown in [18] that the ortho-derivative of a quadratic APN function has algebraic degree at most n2n-2. For any function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, it is well-known that $F$ has algebraic degree $n$ if and only if $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)\neq 0$. Therefore, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic APN, then $\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0$, implying that all component functions of $\pi_{F}$ have even Hamming weight as for any $v\neq 0$, we have $\operatorname{wt}(v\cdot\pi_{F})\equiv\sum_{x\in\mathbb{F}_{2}^{n}}v\cdot\pi_{F}(x)\pmod{2}=v\cdot\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0.$ This allows us to prove that the exclude multiplicities of the graph of an APN quadratic function are always odd.
@@
Assume $n\geq 3$. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Hence, $|\mathcal{B}(F)\cap H|$ is even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$.
@@
If $n$ is odd, then $F$ is AB and $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-2}{6}$ for all $(a,b)\notin\mathcal{G}_{F}$. So, assume $n$ is even, and without loss of generality, assume $F(0)=0$. As mentioned above, the algebraic degree of $\pi_{F}$ is at most n2n-2, implying every component function of $\pi_{F}$ has even Hamming weight. Therefore, by eq. 9, for any $b\neq 0$, we know that $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F}))$ is odd. Thus, by applying Proposition 4.4, we know that any exclude point of $\mathcal{G}_{F}$ has odd multiplicity. The parity of $|\mathcal{B}(F)\cap H|$ must then be even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$, see Proposition 4.14. ∎
@@
As a corollary of Proposition 4.25, every quadratic APN function has a maximal Sidon set as its graph, and as previously mentioned, this is already known since all quadratic functions are plateaued. Nonetheless, it is interesting that Proposition 4.25 now provides a fourth method (the other three being the work of [6], the usage of the D-property in [12], and Theorem 4.13) for proving no quadratic APN function has a non-maximal Sidon set as its graph.

### v2

@@ 4.2 Properties of the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is plateaued APN
Our lower bound in Theorem 4.13 is a dramatic increase over the previously known $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 1$ (for $n\geq 3$) when $F$ is plateaued APN. It would be interesting to further improve this lower bound or find an example of a family of plateaued APN functions for which it is tight. The lower bound on the distance from plateaued APN functions to other APN functions immediately follows.
@@
When $n$ is odd, any plateaued APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is AB, and so we know by a result of [36] that $\mathcal{G}_{F}$ is a maximal Sidon set with all points outside of the graph having exclude multiplicity $\frac{2^{n}-2}{6}$ which is odd (see Section 4.1). However, when $n$ is even, the exclude multiplicities of $\mathcal{G}_{F}$ are much more unclear. In the following proposition, we describe the parity of the exclude multiplicities of $\mathcal{G}_{F}$ for $F$ plateaued APN in terms of the Fourier transform of the indicator function $1_{\mathcal{B}({F})}\colon\mathbb{F}_{2}^{n}\to\left\{0,1\right\}$ of $\mathcal{B}({F})$}. This then tells us that the exclude multiplicities of the graph of any plateaued function are always odd (for $n\geq 3$ since $\mathcal{G}_{F}$ is not a maximal Sidon set if $n\leq 2$).
@@
Assume $n\geq 3$. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function, and let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd.
@@
If $n$ is odd, then $F$ is AB and $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd. So, assume $n$ is even. We claim that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(F(a)+b)\equiv 2\mod 4$. For all $v\in\mathbb{F}_{2}^{n}$, let $\lambda_{v}$ denote the amplitude of $v\cdot F$, and let $m_{v}$ be the non-negative integer such that $\lambda_{v}^{2}=2^{n+m_{v}}$. Note that $m_{v}$ is even and $0\leq m_{v}\leq n$ for all $v\in\mathbb{F}_{2}^{n}$. Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and let $c=F(a)+b$. By Proposition 4.7, we have
@@
So $6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\equiv\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\mod 12$. Note that $\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 6\mod 12$ if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ and $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 0\mod 3$. However, we already know that $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)=0$ as $c\neq 0$. So $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$, and this proves our claim.
Now, for $i\in\left\{0,1\right\}$, let $B_{i}=|\left\{v\in{\mathcal{B}({F})}:v\cdot c=i\right\}|$. Then $\widehat{1_{\mathcal{B}({F})}}(c)=|B_{0}|-|B_{1}|=|{\mathcal{B}({F})}|-2|B_{1}|$. It was shown in [31, Proposition 6.5] that the number of bent components of any plateaued function over an even dimension is $2\mod 4$, and so $|{\mathcal{B}({F})}|\equiv 2\mod 4$. Therefore, $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2-2|B_{1}|\mod 4$. So $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ if and only if $|B_{1}|$ is even. Since the parity of $|B_{0}|$ and $|B_{1}|$ is the same, it follows that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even. Recently, it was shown in [1, Proposition 3.1] that for any plateaued APN function with $n$ even, the intersection of $\mathbb{F}_{2}^{n}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$ with any subspace of dimension at least $\frac{n}{2}+1$ is odd. In particular, this implies that $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even. ∎
In Section 4.4, we provide another proof of the above proposition for quadratic APN functions via the ortho-derivative.
We now express the exclude multiplicities of the graph of a plateaued APN function in terms of $\widehat{W_{F}^{4}}$.
@@
$=|\left\{(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(y)+F(z)+F(x+y+z)=c\right\}|$
@@
From the above result, we can also express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of the derivatives of $\gamma_{F}$. In Section 7, we will see why this is relevant and how it is connected to the existence of nontrivial linear structures of $\gamma_{F}$. In [15, Section 6], it was shown that for any APN function $F$, we have $W_{F}^{2}=2^{2n}\delta_{(0,0)}-W_{\gamma_{F}}+2^{n}$, (where $\delta_{(0,0)}$ takes value $1$ at $(0,0)$ and 0 elsewhere). So
@@
Throughout this subsection, we assume that $n$ is even. We say a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is 33-to-11 if every point in the image of $F$ has a preimage of size $3$ except a single point. Note that all 3-to-1 plateaued functions are necessarily APN (see [30, Corollary 8]).
@@
Moreover, [8, Corollary 4] describes the multiplicities of the nonzero elements of the multiset $\left\{F(x)+F(y)+F(x+y):x,y\in\mathbb{F}_{2}^{n}\right\}$ when $F$ is quadratic 33-to-11, and indeed, this can be easily generalized to all plateaued 3-to-1 APN functions.
@@
These two results above (in terms of quadratic 3-to-1 functions) and [6, Proposition 5] (which states that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$ when $F$ is quadratic) were used in [8] to determine the exact exclude multiplicities of $\mathcal{G}_{F}$ and their frequencies in the case of $F$ being a quadratic 33-to-11 function. This can be done because any plateaued function $F$ with all component functions unbalanced satisfies
@@
for all $x,w\in\mathbb{F}_{2}^{n}$ (see [13, Theorem 2]). Therefore, if $F$ is plateaued APN with all unbalanced components and $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, then
@@
Without loss of generality, assume $F(0)=0$. By Theorem 4.17, $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$. The result then immediately follows by applying Proposition 4.4 and Corollary 4.18 (cf. [8]) and using the same proof as [8, Corollary 4]. ∎
@@ 4.4 Quadratic APN functions
$\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F}))$ (9)
with the second equality following from Proposition 4.15.
@@
Assume $n$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$. For any $b\neq 0$, the weight of the component function $b\cdot\pi_{F}$ of the ortho-derivative of $F$ satisfies $\operatorname{wt}(b\cdot\pi_{F})\leq 2^{n}+2-3\cdot 2^{\frac{n}{2}-1}$.
@@
It was also shown in [19] that the ortho-derivative of a quadratic APN function has algebraic degree at most n2n-2. For any function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, it is well-known that $F$ has algebraic degree $n$ if and only if $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)\neq 0$. Therefore, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic APN, then $\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0$, implying that all component functions of $\pi_{F}$ have even Hamming weight as for any $v\neq 0$, we have $\operatorname{wt}(v\cdot\pi_{F})\equiv\sum_{x\in\mathbb{F}_{2}^{n}}v\cdot\pi_{F}(x)\pmod{2}=v\cdot\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0.$ This allows us to prove that the exclude multiplicities of the graph of an quadratic APN function are always odd in a different manner than Proposition 4.14.
@@
Assume $n\geq 3$. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Equivalently, $|\mathcal{B}(F)\cap H|$ is even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$.
@@
If $n$ is odd, then $F$ is AB and $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-2}{6}$ for all $(a,b)\notin\mathcal{G}_{F}$. So, assume $n$ is even, and without loss of generality, assume $F(0)=0$. As mentioned above, the algebraic degree of $\pi_{F}$ is at most n2n-2, implying every component function of $\pi_{F}$ has even Hamming weight. Therefore, by eq. 9, for any $b\neq 0$, we know that $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F}))$ is odd. Thus, by applying Proposition 4.4, we know that any exclude point of $\mathcal{G}_{F}$ has odd multiplicity. ∎
@@
As a corollary of Proposition 4.24, every quadratic APN function has a maximal Sidon set as its graph, and as previously mentioned, this is already known since all quadratic functions are plateaued. Nonetheless, it is interesting that Proposition 4.24 now provides a fourth method for proving no quadratic APN function (for $n\geq 3$) has a non-maximal Sidon set as its graph (the other three methods being the work of [7], the usage of the D-property in [16], and Theorem 4.13).